### v1

# Bell–Plesset Effects on Rayleigh Taylor Instability of Three Dimensional Spherical Geometry

Abstract
We develop a weakly nonlinear, multi-mode theory for the Rayleigh–Taylor instability (RTI) on a time-varying spherical interface, fully incorporating mode couplings and the Bell–Plesset (BP) effects arising from interface convergence. Our model extends prior analyses, which have been largely restricted to static backgrounds, 2D cylindrical geometries, or single-mode initial condition. We present a framework capable of evolving arbitrary, fully three-dimensional initial perturbations on a dynamic background. At the first order, mode amplitudes respond to the time-varying interface acceleration with an exponential-like growth, in qualitative agreement with classic static results. At second order, nonlinear mode coupling reveals a powerful selection rule: energy is preferentially channeled into axisymmetric ($m=0$) modes. We find that the BP effects dramatically amplify the instability growth by a few orders of magnitude, with this amplification being even more significant for second order couplings. Despite this strong channeling, the second order amplitudes remain small relative to the first order, validating the perturbative approach. These findings offer new physical insights into time-dependent interface instabilities relevant to applications such as astrophysical shell collapse and inertial confinement fusion, highlighting the uniquely dominant role of axisymmetric modes in BP-driven convergent flows.
I Introduction
The Rayleigh Taylor instability (RTI) [RTI1, RTI2] is a fundamental fluid instability that arises at the interface between two fluids of different densities when the lighter fluid supports the heavier one. RTI is a critical mechanism in a wide range of physical systems, from astrophysical phenomena such as supernova explosions and nebula formation [SNe1, SNe2, Nebula1, Nebula2] to technological applications like inertial confinement fusion (ICF) [Hurricane1, Hurricane2, Hurricane3], and meteorological flows [meteorology1, meteorology2, meteorology3]. While the linear stage of RTI is well understood, its dynamics become substantially more complex in the nonlinear regime. Weakly nonlinear (WN) analyses and specialized experiments under idealized conditions have been instrumental in revealing the rich evolutionary behaviors that characterize this stage [RTI_WN1, RTI_WN2], including characteristic structures of penetrating bubbles and spikes [bubble_spike1, bubble_spike2, bubble_spike3].
Analytic investigations of RTI typically employ perturbative expansions to describe the evolution of the fluid interface [RTI_2d_MultiMode, perturbation1, perturbation2, perturbation3, RTI_3d_spherical2]. In a static spherical geometry, the interface radius and the velocity potentials of the two fluids are expanded in spherical harmonics $Y_{l,m}$. The first-order solution shows that each mode grows exponentially and independently, with a characteristic growth rate
$\gamma_{l}=\sqrt{\frac{l(l+1)(\rho_{\rm ex}-\rho_{\rm in})\,g}{\left[(l+1)\rho_{\rm in}+l\rho_{\rm ex}\right]R}}.$ (1)
At higher orders, nonlinear mode coupling emerges, and the strongest growth typically occurs for azimuthal indices near $m\simeq(l+1)/2$ [RTI_3d_spherical1, RTI_3d_spherical2]. However, previous studies compute the higher order evolution only for single-mode initial perturbations and therefore cannot describe the RTI dynamics for an arbitrary interface shape. Moreover, these analyses are fundamentally altered in realistic settings where RTI develops on a dynamic background and Bell-Plesset (BP) effects become significant.
The complexity of RTI is amplified in converging geometries, such as those found in imploding ICF capsules or the collapsing cores of massive stars. In these systems, the time-dependent curvature and background compression introduce the BP effects, which profoundly alters the instability’s growth rates and nonlinear behaviors [BPe1, BPe2, BPe3]. Previous studies of cylindrical implosions (equivalently 2D spherical geometry) have shown that under such conditions, mode coupling emerges beyond the first order approximation, fundamentally altering the RTI dynamics [Zhang2024]. Moreover, BP effects introduce considerable complexities into the evolution equations, precluding closed-form analytic solutions. Although extensive research has explored BP-modified RTI in cylindrical geometries [BPe4, BPe5, BPe6, BPe7], a comprehensive understanding of these effects on a fully three-dimensional spherical interface remains a formidable challenge, primarily due to the high dimensionality of the problem.
Motivated by these considerations, this paper presents a theoretical investigation of RTI development in a fully three-dimensional spherical convergent flow. We carry the perturbation expansion to second order to systematically examine the interplay between Bell–Plesset effects and multi-mode interactions. We organize this paper as follows. Section II introduces our theoretical framework and fundamental equations. Section III describes the setup and numerical methods to solve the equations. Section IV reports the evolution of first order (IV.1) and second order (IV.2) interface perturbations for single-mode and “bubble”-like initial conditions (IV.3), and compares results with and without BP effects (IV.4). Finally, Section V summarizes our conclusions and discusses the broader implications.
II theoretical framework
Our analysis is conducted in a spherical coordinate system, with $r$, $\theta$, and $\varphi$ denoting the radial, poloidal, and azimuthal coordinates, respectively. The interface between the two fluids is described by the perturbed radius
$r_{\text{sf}}(\theta,\varphi,t)=R(t)+\eta(\theta,\varphi,t),$ (2)
where $R(t)$ is the radius of the unperturbed interface and $\eta$ is the perturbations. The velocity field is represented by the potential $\Phi^{i}=\psi+\phi^{i}$, and $i=\mathrm{ex}\text{ or }\mathrm{in}$ denotes the exterior and interior components, $\psi$ describes the background flow (depending only on $r$ and $t$), and $\phi^{i}$ are perturbative contributions. We consider ideal, incompressible fluids with zero viscosity. Combining the equations of conservations of mass and momentum, we obtain
$\nabla^{2}\phi^{i}=0,$ (3)
$\begin{split}\frac{\partial\eta}{\partial t}&+\frac{1}{r^{2}\sin^{2}\theta}\frac{\partial\eta}{\partial\varphi}\frac{\partial\phi^{i}}{\partial\varphi}+\frac{1}{r^{2}}\frac{\partial\eta}{\partial\theta}\frac{\partial\phi^{i}}{\partial\theta}\\ &-\frac{\partial\phi^{i}}{\partial r}+\left(\dot{R}-\frac{\partial\psi}{\partial r}\right)=0\quad\text{at }r=R+\eta,\end{split}$ (4)
$\begin{split}\biggl\{&\frac{1}{2}\bigg[\frac{1}{r^{2}\sin^{2}\theta}\left(\frac{\partial\Phi^{i}}{\partial\varphi}\right)^{2}+\frac{1}{r^{2}}\left(\frac{\partial\Phi^{i}}{\partial\theta}\right)^{2}+\left(\frac{\partial\Phi^{i}}{\partial r}\right)^{2}\bigg]\\ &+\frac{\partial\Phi^{i}}{\partial t}+gr\biggr\}\Bigg|^{i=\text{ex}}_{i=\text{in}}=f(t)\quad\text{at }r=R+\eta.\end{split}$ (5)
Here, $f(t)$ is an arbitrary function of time, which we henceforth take to be zero, and $g=\ddot{R}$ denotes the acceleration of the interface. Notably, $g$ is distinct from $\ddot{R}$ in that it represents the external force driving the instability. In the literature, $g$ often appears as gravitational acceleration; in our setting, however, external driving forces dominate while the gravity can be neglected.
To determine $\psi$, we define the compression rate
$\gamma_{\rho}=\frac{\dot{\rho}}{\rho}=-\frac{3}{r^{3}-R^{3}}(r^{2}\dot{r}-R^{2}\dot{R}),$ (6)
where $\rho$ is the density at $R$, and $r$ is an arbitrary radius near $R$, assuming homogeneity within this thin layer. We then obtain
$\frac{\partial\psi}{\partial r}=\dot{r}=\frac{R^{2}\dot{R}}{r^{2}}+\frac{\gamma_{\rho}}{3}\left(\frac{R^{3}}{r^{2}}-r\right).$ (7)
We then expand these functions in spherical harmonics:
$\eta(\theta,\varphi,t)=\sum_{l,m}a_{l,m}(t)Y_{l}^{m}(\theta,\varphi),$ (8)
$\phi^{\rm{ex}}(\theta,\varphi,t)=\sum_{l,m}b^{\rm{ex}}_{l,m}(t)\left(\frac{r}{R}\right)^{-(l+1)}Y_{l}^{m}(\theta,\varphi),$ (9)
$\phi^{\rm{in}}(\theta,\varphi,t)=\sum_{l,m}b^{\rm{in}}_{l,m}(t)\left(\frac{r}{R}\right)^{l}Y_{l}^{m}(\theta,\varphi),$ (10)
where these forms automatically satisfy Eq. (3) and the boundary conditions $\vec{v}|_{r=0}$ is finite and $\vec{v}|_{r\to\infty}=0$. The amplitudes are then expanded in a formal small parameter $\epsilon$:
$a_{l,m}=\epsilon a^{(1)}_{l,m}+\epsilon^{2}a^{(2)}_{l,m}$$\cdots$, (11)
$b^{\rm{in}}_{l,m}=\epsilon b^{\text{in}(1)}_{l,m}+\epsilon^{2}b^{\text{in}(2)}_{l,m}$$\cdots$, (12)
$b^{\rm{ex}}_{l,m}=\epsilon b^{\text{ex}(1)}_{l,m}+\epsilon^{2}b^{\text{ex}(2)}_{l,m}$$\cdots$. (13)
Substituting the above expansions into Eq. (4) and Eq. (5) then yields the full coupled system (explicit expressions are given in Appendix A). We take $\epsilon=1$ to better compare first and second order amplitudes. Solving these equations order by order, we botain the amplitudes for arbitrary modes.
We discuss the properties of first and second order equations in the following of this section.
II.1 First Order Equation
At first order, $b^{\text{in}(1)}_{l,m}$ and $b^{\text{ex}(1)}_{l,m}$ are related to $a^{(1)}_{l,m}$ as
$b^{\text{in}(1)}_{l,m}=\frac{R\dot{a}^{(1)}_{l,m}+(2\dot{R}+R\gamma_{\rho})a^{(1)}_{l,m}}{l},$ (14)
$b^{\text{ex}(1)}_{l,m}=-\frac{R\dot{a}^{(1)}_{l,m}+(2\dot{R}+R\gamma_{\rho})a^{(1)}_{l,m}}{l+1},$ (15)
Substituting into the momentum conservation equation yields
[…]
The first-order evolution is independent of azimuthal number $m$. In the static limit where there’s no BP effects ($R=\text{const}$, $\gamma_{\rho}=0$), the reduced growth factor is same with Eq.(1), thus validating our derivations in none-BP limit.
To better understand this result, we simplify Eq. (LABEL:eq:first_order_eq) by defining
$\gamma_{R}=\frac{\dot{R}}{R},\quad u=\frac{2l(l+1)At}{1+2l+At},$ (17)
where $At=\frac{\rho^{\mathrm{in}}-\rho^{\mathrm{ex}}}{\rho^{\mathrm{in}}+\rho^{\mathrm{ex}}}$ is the Atwood number [Atwood1, Atwood2]. The evolution equation ((LABEL:eq:first_order_eq)) then becomes
$\ddot{a}+(\gamma_{\rho}+3\gamma_{R})\dot{a}+\left[\gamma_{\rho}\gamma_{R}+\dot{\gamma}_{\rho}+\frac{ug}{R}+(u+2)(\dot{\gamma}_{R}+\gamma_{R}^{2})\right]a=0.$ (18)
Let $x(t)$ and $y(t)$ denote the coefficients of $\dot{a}$ and $a$
$x(t)=\gamma_{\rho}+3\gamma_{R},$ (19)
$y(t)=\gamma_{\rho}\gamma_{R}+\dot{\gamma}_{\rho}+\frac{ug}{R}+(u+2)(\dot{\gamma}_{R}+\gamma_{R}^{2}),$ (20)
and define
$a(t)=b(t)\exp\left[-\frac{1}{2}\int x(t)\mathrm{d}t\right]=\frac{C_{0}}{\sqrt{\rho R^{3}}}b(t),$ (21)
where $C_{0}$ is the integral constant. The equation then simplifies to
$\ddot{b}+z(t)b=0,$ (22)
where $z(t)=y(t)-\frac{1}{4}\dot{x}(t)-\frac{1}{2}x^{2}(t)$.
Generally, this equation lacks an analytic solution. The sign of $z(t)$ determines instability: $z<0$ produces exponential-like amplification, while $z>0$ results in sinusoidal fluctuations and thus suppresses the instability. Since it exhibits scale invariance with respect to the initial amplitude and independent from the azimuthal index $m$, we solve for a normalized growth factor for each mode number $l$, defined as
$g_{l}(t)=\frac{b_{l}(t)}{b_{l}(t_{0})}=\frac{a^{(1)}_{l,m}(t)}{a^{(1)}_{l,m}(t_{0})},$ (23)
which is identical for all $m$ at a given $l$. This growth factor obeys the same differential form as Eq. (22), which is subjected to the initial conditions of a quiescent interface
$g_{l}(t_{0})=1,\quad\dot{g}_{l}(t_{0})=0.$ (24)
II.2 Second Order Equations
The second-order equations are considerably complex and their full derivations are presented in Appendix B.
The mass conservation equations for the internal component is given by
$\begin{split}&\frac{2}{3}\left[\dot{a}^{(2)}_{l,m}+\left(2\gamma_{R}+\gamma_{\rho}\right)a^{(2)}_{l,m}-b^{\text{in}(2)}_{l,m}\frac{l}{R}\right]+\sum_{l_{1},m_{1}}\bigg[\dot{a}^{(2)}_{l_{1},m_{1}}+\left(2\gamma_{R}+\gamma_{\rho}\right)a^{(2)}_{l_{1},m_{1}}-b^{\text{in}(2)}_{l_{1},m_{1}}\frac{l_{1}}{R}\bigg]A_{l,m,l_{1},m_{1}}+\frac{1}{R^{2}}\sum_{l_{1},m_{1},l_{2},m_{2}}\bigg\{a^{(1)}_{l_{1},m_{1}}b^{\text{in}(1)}_{l_{2},m_{2}}\\ &\quad\times\Theta_{l,m,l_{1},m_{1},l_{2},m_{2}}-m_{1}m_{2}a^{(1)}_{l_{1},m_{1}}b^{\text{in}(1)}_{l_{2},m_{2}}B_{l,m,l_{1},m_{1},l_{2},m_{2}}-\big[3\dot{R}+R\gamma_{\rho}+l_{2}(l_{2}-1)a^{(1)}_{l_{1},m_{1}}b^{\text{in}(1)}_{l_{2},m_{2}}\big]H_{l,m,l_{1},m_{1},l_{2},m_{2}}\bigg\}=0,\end{split}$ (25)
and for the external component we have
$\begin{split}\frac{2}{3}&\left[\dot{a}^{(2)}_{l,m}+\left(2\gamma_{R}+\gamma_{\rho}\right)a^{(2)}_{l,m}+b^{\text{ex}(2)}_{l,m}\frac{l+1}{R}\right]+\sum_{l_{1},m_{1}}\left[\dot{a}^{(2)}_{l_{1},m_{1}}+\left(2\gamma_{R}+\gamma_{\rho}\right)a^{(2)}_{l_{1},m_{1}}+b^{\text{ex}(2)}_{l_{1},m_{1}}\frac{l_{1}+1}{R}\right]A_{l,m,l_{1},m_{1}}\\ +&\frac{1}{R^{2}}\sum_{l_{1},m_{1},l_{2},m_{2}}\bigg\{a^{(1)}_{l_{1},m_{1}}b^{\text{ex}(1)}_{l_{2},m_{2}}\Theta_{l,m,l_{1},m_{1},l_{2},m_{2}}-m_{1}m_{2}a^{(1)}_{l_{1},m_{1}}b^{\text{ex}(1)}_{l_{2},m_{2}}B_{l,m,l_{1},m_{1},l_{2},m_{2}}-\\ &\big[3\dot{R}+R\gamma_{\rho}+(l_{2}+1)(l_{2}+2)a^{(1)}_{l_{1},m_{1}}b^{\text{ex}(1)}_{l_{2},m_{2}}\big]H_{l,m,l_{1},m_{1},l_{2},m_{2}}\bigg\}=0,\end{split}$ (26)
where $A_{l,m,l_{1},m_{1}}$, $B_{l,m,l_{1},m_{1},l_{2},m_{2}}$, $H_{l,m,l_{1},m_{1},l_{2},m_{2}}$, and $\Theta_{l,m,l_{1},m_{1},l_{2},m_{2}}$ are quantities defined in Appendix B.
The momentum conservation equation is expressed as
$\begin{split}\rho^{\mathrm{in}}&\Bigg\{\frac{2}{3}\left[(\ddot{R}+g){a}^{(2)}_{l,m}+\dot{b}^{\text{in}(2)}_{l,m}\right]+\sum_{l_{1},m_{1}}\bigg[(\ddot{R}+g){a}^{(2)}_{l_{1},m_{1}}+\dot{b}^{\text{in}(2)}_{l_{1},m_{1}}\bigg]A_{l,m,l_{1},m_{1}}+\sum_{l_{1},m_{1},l_{2},m_{2}}\Bigg[\frac{l_{2}}{R}a^{(1)}_{l_{1},m_{1}}\dot{b}^{\text{in}(1)}_{l_{2},m_{2}}-l_{2}^{2}\frac{\gamma_{R}}{R}a^{(1)}_{l_{1},m_{1}}b^{\text{in}(1)}_{l_{2},m_{2}}\\ &+\frac{1}{R^{2}}\bigg(\dot{R}^{2}-R\ddot{R}-\frac{\dot{\gamma}_{\rho}R^{2}}{2}\bigg)a^{(1)}_{l_{1},m_{1}}a^{(1)}_{l_{2},m_{2}}+\frac{1}{R}\Big(l_{2}(l_{2}-1)\gamma_{R}-l_{2}(2\gamma_{R}+\gamma_{\rho})\Big)a^{(1)}_{l_{1},m_{1}}b^{\text{in}(1)}_{l_{2},m_{2}}+\frac{(2\gamma_{R}+\gamma_{\rho})^{2}}{2}a^{(1)}_{l_{1},m_{1}}a^{(1)}_{l_{2},m_{2}}\\ &+\frac{l_{1}l_{2}}{2R^{2}}b^{\text{in}(1)}_{l_{1},m_{1}}b^{\text{in}(1)}_{l_{2},m_{2}}\Bigg]H_{l,m,l_{1},m_{1},l_{2},m_{2}}+\frac{1}{2R^{2}}\sum_{l_{1},m_{1},l_{2},m_{2}}\bigg(b^{\text{in}(1)}_{l_{1},m_{1}}b^{\text{in}(1)}_{l_{2},m_{2}}\Theta_{l,m,l_{1},m_{1},l_{2},m_{2}}-m_{1}m_{2}b^{\text{in}(1)}_{l_{1},m_{1}}b^{\text{in}(1)}_{l_{2},m_{2}}B_{l,m,l_{1},m_{1},l_{2},m_{2}}\bigg)\Bigg\}\\ &\quad\quad=\\ \rho^{\mathrm{ex}}&\Bigg\{\frac{2}{3}\left[(\ddot{R}+g){a}^{(2)}_{l,m}+\dot{b}^{\text{ex}(2)}_{l,m}\right]+\sum_{l_{1},m_{1}}\bigg[(\ddot{R}+g){a}^{(2)}_{l_{1},m_{1}}+\dot{b}^{\text{ex}(2)}_{l_{1},m_{1}}\bigg]A_{l,m,l_{1},m_{1}}+\sum_{l_{1},m_{1},l_{2},m_{2}}\Bigg[-(l_{2}+1)^{2}\frac{\gamma_{R}}{R}a^{(1)}_{l_{1},m_{1}}b^{\text{ex}(1)}_{l_{2},m_{2}}\\ &-\frac{l_{2}+1}{R}a^{(1)}_{l_{1},m_{1}}\dot{b}^{\text{ex}(1)}_{l_{2},m_{2}}+\frac{1}{R^{2}}\bigg(\dot{R}^{2}-R\ddot{R}-\frac{\dot{\gamma}_{\rho}R^{2}}{2}\bigg)a^{(1)}_{l_{1},m_{1}}a^{(1)}_{l_{2},m_{2}}+\frac{1}{R}\Big((l_{2}+1)(l_{2}+2)\gamma_{R}+(l_{2}+1)(2\gamma_{R}+\gamma_{\rho})\Big)a^{(1)}_{l_{1},m_{1}}b^{\text{ex}(1)}_{l_{2},m_{2}}\\ &+\frac{(2\gamma_{R}+\gamma_{\rho})^{2}}{2}a^{(1)}_{l_{1},m_{1}}a^{(1)}_{l_{2},m_{2}}+\frac{(l_{1}+1)(l_{2}+1)}{2R^{2}}b^{\text{ex}(1)}_{l_{1},m_{1}}b^{\text{ex}(1)}_{l_{2},m_{2}}\Bigg]H_{l,m,l_{1},m_{1},l_{2},m_{2}}\\ &+\frac{1}{2R^{2}}\sum_{l_{1},m_{1},l_{2},m_{2}}\bigg(b^{\text{ex}(1)}_{l_{1},m_{1}}b^{\text{ex}(1)}_{l_{2},m_{2}}\Theta_{l,m,l_{1},m_{1},l_{2},m_{2}}-m_{1}m_{2}b^{\text{ex}(1)}_{l_{1},m_{1}}b^{\text{ex}(1)}_{l_{2},m_{2}}B_{l,m,l_{1},m_{1},l_{2},m_{2}}\bigg)\Bigg\},\end{split}$ (27)
To make the computation tractable, we truncate the spherical harmonic expansion at a maximum mode number of $l_{\max}$. This truncation yields a system of $N=\sum_{l=1}^{l_{\max}}(2l+1)=(l_{\max}+1)^{2}-1$ distinct modes. This system can then be cast into a compact matrix form:
$U\left(\dot{A}+(2\gamma_{R}+\gamma_{\rho})A\right)+V_{i}B_{i}+Q_{i}=0,$ (28)
$\biggl\{\rho_{i}\left(U(\dot{B}_{i}+\ddot{R}A+gA)+P_{i}\right)\biggr\}\Bigg|^{i=\mathrm{ex}}_{i=\mathrm{in}}=0.$ (29)
Here, $A$, $B_{\text{in}}$, and $B_{\text{ex}}$ are $N\times 1$ column vectors containing the amplitudes $a^{(2)}_{l,m}$, $b^{\text{in}(2)}_{l,m}$, and $b^{\text{ex}(2)}_{l,m}$, respectively. The elements are ordered using the index mapping
$k=l^{2}+l+m.$ (30)
The quantities $U$, $V_{i}$ are $N\times N$ coefficient matrices, and $P_{i}$, $Q_{i}$ are $N\times 1$ vectors, with the same index mapping. Their explicit forms are provided in Appendix C. Notably, the matrices $V_{i}$, $P_{i}$, and $Q_{i}$ are time-dependent and contain the first-order amplitudes, whereas $U$ is a constant matrix.
Eliminate $\dot{A}$ from Eq. (28), we obtain
$B_{\text{ex}}=V_{\text{ex}}^{-1}(V_{\text{in}}B_{\text{in}}+Q_{\text{in}}-Q_{\text{ex}}).$ (31)
Plugging Eq. (31) into Eq. (29) yields a closed form evolution equation for $B_{\text{in}}$:
$\begin{split}&(\rho_{\text{in}}U-\rho_{\text{ex}}UV_{\text{ex}}^{-1}V_{\text{in}})\dot{B}_{\text{in}}-\rho_{\text{ex}}UV_{\text{ex}}^{-1}\dot{V}_{\text{in}}B_{\text{in}}+\rho_{\text{in}}P_{\text{in}}\\ &-\rho_{\text{ex}}P_{\text{ex}}+(\rho_{\text{in}}-\rho_{\text{ex}})(\ddot{R}+g)UA-\rho_{\text{ex}}UV_{\text{ex}}^{-1}(\dot{Q}_{\text{in}}-\dot{Q}_{\text{ex}})\\ &-\rho_{\text{ex}}U\dot{V}_{\text{ex}}^{-1}(V_{\text{in}}B_{\text{in}}+Q_{\text{in}}-Q_{\text{ex}})=0.\end{split}$ (32)
We integrate this system with initial conditions $A(t_{0})=B_{\text{in}}(t_{0})=0$ and $\dot{A}(t_{0})=\dot{B}_{\text{in}}(t_{0})=0$, which correspond to zero initial second-order perturbations. The evolution of $A(t)$ and $B_{\text{in}}(t)$ follows Eq. (28) and (32).
III Background Setup
Before integrating the system, we must prescribe the time dependence of $R(t)$, $\rho^{i}(t)$ and $\gamma_{\rho}(t)$. In realistic settings these arise from detailed simulations or experiments, but to explore general behavior we adopt an analytic $R(t)$ to mimic implosion, impose the relation $\rho^{i}R^{3}=\mathrm{const}$ which implies $\gamma_{\rho}=-3\frac{\dot{R}}{R}$, and give the initial condition $\rho^{\mathrm{ex}}=5\rho^{\mathrm{in}}$, with Atwood number corresponds to $At=-\frac{2}{3}$. We note here that the same formalism would apply equally well to an explosion or Atwood number as an arbitrary function of time.
Figure 1: Demonstration of the implosion prescription: the blue curve shows $R/R_{f}$ as a function of $\gamma t$ (left axis), while the red curve shows $\dot{R}/(\gamma R_{f})$ vs. $\gamma t$ (right axis).
A realistic implosion rate tends to rise, then decline toward stagnation. We emulate such implosion process by constructing
$R(t)=R_{f}+(R_{0}-R_{f})\frac{1+e^{-\lambda_{0}}}{1+e^{\gamma t-\lambda_{0}}},$ (33)
where $R_{0}$ and $R_{f}$ are the initial and final radii, $\lambda_{0}$ controls the start stage, and $\gamma$ controls the collapse speed. For our numerical studies we choose $R_{0}=5\,R_{f}$, $\lambda_{0}=5$. Figure 1 shows the resulting interface radius $R(t)$ (normalized to $R_{f}$) and implosion rate $\dot{R}(t)$ (normalized to $\gamma R_{f}$) from $\gamma t=0$ to $\gamma t=10$.
Figure 2: Time evolution of $\gamma^{2}z_{l}(t)$ for modes $l=1,5,10$ (solid lines), and the scaled acceleration $\gamma^{2}\,\ddot{R}/R_{f}$ (dashed line). The similar shapes highlight the tight relation between $z_{l}$ and the interface acceleration.
In Figure 2 we plot $\gamma^{2}z_{l}(t)$ for $l=1,5,10$ alongside the scaled acceleration $\gamma^{2}\ddot{R}/R_{f}$ (recall from Eq. 22 that the normalized amplitude $b_{l}$ satisfies $\ddot{b}+z_{l}(t)\,b=0$). Notably, the temporal profiles align closely, showing how the instantaneous interface acceleration directly influences $z_{l}$. As $l$ increases, the amplitude of $\gamma^{2}z_{l}$ also increases, consistent with the built-in scaling of $z_{l}$. When $\ddot{R}=0$, $\gamma^{2}z_{l}$ crosses zero (unlike static sphere cases where gravitational acceleration enforces a nonzero value). Early in the implosion, $\gamma^{2}z_{l}$ is positive, implying oscillatory behavior. However, during the deceleration stage it becomes negative, driving exponential-like growth. This suggests that in a converging interface, higher-$l$ modes experience more intense early oscillations and more rapid late growth, in qualitative alignment with static background expectations.
During the early stage of the implosion ($\gamma t<5$), the interface acceleration is directed outward, leading to $z_{l}<0$ and suppressing the Rayleigh–Taylor instability (RTI). In this regime, the perturbation amplitudes merely oscillate rather than grow, rendering it unimportant for studying RTI. Therefore, we initiate the perturbation evolution at $\gamma t=5$.
For the full evolution, including both first- and second order amplitudes, the physical requirement that the interface perturbation $\eta(\theta,\varphi,t)$ must be real-valued imposes the condition
$a_{l,-m}(t)=(-1)^{m}\,a^{*}_{l,m}(t)\;.$
Thus, initializing a single complex spherical harmonic mode would yield a non-physical complex interface shape. To ensure real-valued perturbations, we consider single-mode static initial conditions with real amplitudes and vanishing initial velocities at $\gamma t_{0}=5$:
$a_{l_{0},m_{0}}(t_{0})=(-1)^{m_{0}}a_{l_{0},-m_{0}}(t_{0})=\,\eta_{0},\quad\dot{a}^{(1)}_{l,m}=0,$
for the representative cases $(l_{0},m_{0})=(3,1),\,(4,0),\,(4,1),\,(4,4)$. The axisymmetric mode ($m=0$) is inherently real, so we set $a_{4,0}(t_{0})=2\,\eta_{0}$ to ensure same amplitude. In all cases, we take $\eta_{0}/R_{f}=0.002$.
To explore more physically realistic RTI behavior, we further initialize the interface with a Gaussian-shaped bubble centered on the north pole:
$\eta(\theta,\varphi,t_{0})=0.004\times\exp(-10\theta^{2}).$ (34)
This configuration mimics a localized deformation of the interface and illustrates how such a perturbation evolves over time.
As discussed in Section II, we truncate the mode at $l_{\text{max}}=16$ to avoid unaffordable computational cost. For most interested low-$l$ modes, $l_{\text{max}}=16$ ensures acceptable numerical convergence (details can be found in Appendix D).
IV Numerical Results
In this section, we present the outcomes of the numerical integration of the first and second order perturbation equations and interpret their physical consequences. We begin with the single mode evolution, highlighting how a dynamic background (i.e., Bell–Plesset effects) modifies first order behavior relative to static spherical geometry. We then turn to multi-mode interactions arising from second order coupling, exploring energy diffusion among modes and its impact on interface growth. After that, we analyze the “bubble”-like structure evolution, studying how multi-mode initial condtion influence our results. Finally, we analyze the influence of the BP effects, highlighting its role in amplifying the instability.
IV.1 First Order Evolution
At the first order, the different spherical harmonic modes evolve independently. Figure 3 presents a heatmap of the first order growth factor, $g_{l}(t)$, for poloidal mode numbers ranging from $l=1$ to $20$. The visualization clearly illustrates a primary characteristic of the first order amplitudes that higher $l$ modes grow faster over the same time interval, consistent with the larger $\gamma_{l}$ (Eq. (1)) found in static background.
Figure 3: Heatmap of $g_{l}(t)$ for $l=1$ to 2020 evolving from $\gamma t=5$ to $10$. Redder regions denote large amplitudes, while blue regions denote small amplitudes.
To illustrate individual mode behavior, Figure 4 plots $g_{l}(t)$ for selected $l=1,2,5,10,20$ in log scale. The amplitudes start with an initial period of slow amplification, followed by a phase of rapid, super-exponential growth, and finally a transition toward slower, linear growth as the driving term $z_{l}(t)\to 0$ at last in our prescribed implosion scenario. While higher-$l$ modes achieve much larger amplitudes, the temporal profile of their growth is qualitatively similar across all modes, in correspondance with the $z_{l}$ shown in Figure 2. Our results show that in a dynamically imploding background, the first order evolution behaves qualitatively similar to static sphere models, though BP effects enhances the instability growth by adding extra terms in Eq. (LABEL:eq:first_order_eq).
Figure 4: Evolution of the growth factor $g_{l}(t)$ for selected modes ($l=1,2,5,10,20$) plotted on a logarithmic scale. The curves illustrate a period of accelerated growth followed by a saturation into a linear growth regime at late times as the interface deceleration ceases.
IV.2 Second Order Evolution
Moving to the second order analysis, we now solve the full coupled system to investigate the nonlinear dynamics, with the four different initial conditions $(l,m)=(3,\pm 1),\,(4,0),\,(4,\pm 1),\,(4,\pm 4)$.
Figure 5 displays the interface morphology for each case at the final time $\gamma t=10$. The surfaces exhibit complex, multi-modal topographies, a stark departure from the simple, single-mode structure of the initial conditions. This complexity is a direct consequence of nonlinear mode coupling, which redistributes energy from the initial modes to a broad spectrum of other modes.
Figure 5: Interface shape at $\gamma t=10$ for four initial conditions: $(l,m)=\text{(a) }(3,\pm 1),\,\text{(b) }(4,0),\,\text{(c) }(4,\pm 1),\,\text{(d) }(4,\pm 4)$. Colors represent the normalized displacement amplitude ($\eta/R_{f}$), where red indicates outward displacement ($\eta>0$) and blue indicates inward displacement ($\eta<0$). The directions of coordinates are marked by the arrows on the bottom left.
A striking feature in Figure 5 is the markedly larger deformation observed for the axisymmetric $(l,m)=(4,0)$ case compared to the others. This observation strongly suggests that axisymmetric perturbations are predisposed to more vigorous growth. The spectral analysis presented next provides a quantitative confirmation of this behavior.
[…]
To explain this phenomenon, we must examine the structure of the coupling coefficients in Eqs. (25)-(27). The $B_{l,m,l_{1},m_{1},l_{2},m_{2}}$, $H_{l,m,l_{1},m_{1},l_{2},m_{2}}$, and $\Theta_{l,m,l_{1},m_{1},l_{2},m_{2}}$ terms act as the source terms that couple two first-order quantities to generate a second-order amplitude. The $A_{l,m,l_{1},m_{1}}$ term governs the subsequent couplings among the second-order amplitudes themselves. As detailed in Appendix B, all these coefficients are derived from integrals over products of several spherical harmonics.
The key to the selection rule lies in the inversion symmetry (parity) of spherical harmonics:
$Y_{l}^{m}(\pi-\theta,\pi+\varphi)=(-1)^{l}Y_{l}^{m}(\theta,\varphi).$ (35)
For an integral over the full sphere of a product of spherical harmonics ($\prod_{i}Y_{l_{i}}^{m_{i}}$) to be non-zero, the integrand must be an even function under this transformation, as the integral of any odd function over this symmetric domain is zero. This imposes a strict parity selection rule: the integral vanishes unless the sum of the polar indices, $\sum l_{i}$, is an even number.
We first apply this rule to the source terms ($B$, $H$, $\Theta$). These coefficients couple two first-order modes ($l_{1}$, $l_{2}$) to generate a second-order mode ($l$). The integral is non-zero only if $l_{1}+l_{2}+l=\text{even}$. In our case, the first-order modes originate from the same initial mode, $l_{0}$, so $l_{1}=l_{2}=l_{0}$. The condition for excitation thus becomes $l_{0}+l_{0}+l=\text{even}$. Since $l_{0}+l_{0}$ is always even, this rigorously forces the resulting second-order mode $l$ to be even.
Figure 6: Spectra of normalized second-order amplitudes $|a^{(2)}_{l,m}|/R_{f}$ at $\gamma t=10$ for four initial conditions. The horizontal axis is the normalized azimuthal index $m/l$. Solid curves (even $l$) and dashed curves (odd $l$) show a powerful selection rule: only the equatorially symmetric (even $l$) modes are excited, and among these, the axisymmetric ($m=0$) modes are dominant.
Once these even-$l$ modes are generated, the $A_{l,m,l_{1},m_{1}}$ coefficient, which governs their mutual interactions, must also obey the same parity rule. This ensures that an even $l$ mode can only couple to another even $l_{1}$ mode, creating a closed system for even-$l$ modes. As a result, energy diffuses exclusively among even-$l$ modes. This is precisely what is observed in Figure 6: only the even-$l$ amplitudes (solid lines) are excited, while all odd-$l$ amplitudes (dashed lines) are zero.
Additionally, a second selection rule must be satisfied for the azimuthal index: $\sum m_{i}=0$. This arises because any un-cancelled factor $e^{i\varphi\sum m_{i}}$ in the integrand integrates to zero over $\varphi\in[0,2\pi]$. The $m=0$ modes are unique in that they can be generated by any self-coupling ($m_{1}=m_{0},m_{2}=-m_{0}$) and participate in the largest number of allowed coupling pathways. This “coupling promiscuity” likely explains their preferential growth and observed dominance in the spectra. This analysis validates our finding that the axisymmetric ($m=0$) modes are the fastest growing.
Notably, our results indicate that the azimuthal mode number $m$ of the initial perturbation plays a crucial role in determining the pattern of the instability, whereas the poloidal mode number $l$ primarily governs the instability growth rate. For example, the interface morphology and spectrum for $(l,m)=(3,\pm 1)$ (Figure 5a, 6a) bear a qualitative resemblance to those of the $(l,m)=(4,\pm 1)$ case (Figure 5c, 6c). Moreover, the three cases with $l=4$ but different $m$ values, namely $(4,0)$, $(4,\pm 1)$, and $(4,\pm 4)$, yield distinctly different outcomes. In contrast, the perturbation growth rates for the three cases with the same $l$ value are comparable, and significantly larger than that of the $l=3$ case.
Figure 7: Evolution of the first order amplitude (dashed line) and the dominant second order amplitudes (solid lines, scaled by 1/5 for visibility) for the four initial conditions: $(l,m)=\text{(a) }(3,\pm 1),\,\text{(b) }(4,0),\,\text{(c) }(4,\pm 1),\,\text{(d) }(4,\pm 4)$. The second order amplitudes show complex growth but remain significantly smaller than the first order mode.
To compare the growth of the orders, Figure 7 illustrates the time evolution of the dominant second order modes ($l=2,6,8,12$, $m=0$) alongside the initial first order amplitude (scaled down by a factor of five). The second order amplitudes display accelerating growth rates that eventually surpass that of the first order. Notably, they remain significantly smaller in magnitude than the first order amplitude throughout the simulation, while the first order amplitude itself remains much smaller than the unperturbed interface radius. This result is crucial, as it validates the applicability of our perturbative method up to $\gamma t=10$, confirming that the system remains within the weakly nonlinear regime and our model provides an accurate assessment of the instability. Furthermore, the plots reveal that the dominant second order amplitudes tend to develop negative signs, whereas other modes like $(2,0)$ exhibit positive values, highlighting the complex phase dynamics that are not present in the first order growth.
Figure 8: Interface evolution for the “bubble” perturbation defined in Eq. (34). (a) The initial perturbation shape, reconstructed from the $l\leq 16$ spectral projection. The amplitudes are scaled 50x for visibility purpose. (b) The evolved interface at $\gamma t=10$.
IV.3 “Bubble”-like Structure Evolution
We examine the RTI growth from a more complex, multi-mode initial condition designed to simulate a localized “bubble” perturbation (Eq. (34)). To construct this state, we first define a desired initial surface perturbation, $\eta(\theta,\varphi,t_{0})$, and then project it onto the spherical harmonic basis to obtain the initial modal amplitudes, $a^{(1)}_{l,m}(t_{0})$ via
$a^{(1)}_{l,m}(t_{0})=\int\eta(\theta,\varphi,t_{0})Y^{*}_{l,m}\mathrm{d}\Omega.$ (36)
Figure 8 shows the initial and final interface shapes. The initial shape, as reconstructed from our $l\leq 16$ spectral expansion, demonstrates that our truncation adequately captures the localized Gaussian perturbation. The subsequent evolution shows that the bubble grows locally, largely retaining its shape while increasing in amplitude.
Figure 9: Spectrum of normalized second order amplitudes $|a^{(2)}_{l,m}|/R_{f}$ for $l=4$ to 1212 at $\gamma t=10$, resulting from the multi-mode “bubble” initial condition. The horizontal axis is the normalized azimuthal index $m/l$. Solid (dashed) lines represent even (odd) $l$.
Figure 9 shows the resulting second order spectral distribution at $\gamma t=10$. Since the initial “bubble” state contains a superposition of both even and odd $l$ modes, the selection rule $l_{1}+l_{2}+l=\text{even}$ now permits the excitation of odd-$l$ second order amplitudes, which are clearly visible in the spectrum. Despite the excitation of this broader range of modes, the most significant finding is that the axisymmetric ($m=0$) modes exhibit an even stronger tendency to dominate the nonlinear growth compared to the single-mode initial conditions. This result powerfully reinforces our previous conclusion regarding the robust and preferential instability of $m=0$ perturbations.
A trend also emerges in Figure 9 that the second order amplitudes systematically decrease as their mode number $l$ increases to larger values. This observation lends physical justification to our spectral truncation at $l_{\max}=16$, as higher-$l$ modes are increasingly suppressed, rendering the influence of the unconverged modes limited by the truncation negligible.
Figure 10: Evolution of the normalized bubble height ($\eta_{\text{peak}}/R_{f}$, blue-square curve) and bubble radius ($R_{b}/R_{f}$, orange-circle curve) for the multi-mode “bubble” case. The radius $R_{b}$ is defined as the location where the amplitude drops to 1/10 of the peak.
To better quantify the instability’s growth, we track two metrics: the bubble’s peak height ($\eta_{\text{peak}}$ at the pole) and its radius $R_{b}$. We define the boundary of the bubble as the area where the perturbation amplitude falls to 1/10 of its peak value, the radius is then the distance between the peak and the boundary. Figure 10 plots the evolution of these two quantities. The bubble height grows in a manner qualitatively similar to the first order amplitudes, which is expected as the perturbation growth is dominated by first order. Concurrently, the bubble radius $R_{b}$ slowly decreases. This combined behavior with increasing height and decreasing radius indicates that the bubble sharpens as the evolution progresses.
IV.4 Influence of BP effects
To isolate the effects of a static interface (i.e., suppressing the Bell–Plesset effects), we perform an additional numerical experiment in which the radius is held constant, $R=R_{0}$, while the driving acceleration retains the same functional form $g=\ddot{R}(t)$ from Eq. (33), for the case $(l,m)=(4,\pm 1)$.
Figure 11: Heatmap of the first order growth factor $g_{l}(t)$ for modes $l=1$ to $20$, evolved from $\gamma t=5$ to 10 under the condition of a fixed interface radius (no BP effects).
Figure 11 presents the heatmap of the first order amplitude growth factors under this static interface scenario. Compared to the corresponding plot in Figure 3, the pattern of mode behaviour is identical, but the amplitude is reduced by roughly two orders of magnitude. This clearly demonstrates that although the qualitative structure of the evolution remains, the presence of Bell–Plesset curvature/convergence effects drastically amplifies the growth rate of the instability.
Figure 12: Spectrum of normalized second order amplitudes $\bigl|a^{(2)}_{l,m}\bigr|/R_{f}$ for $l=4$ to 12 at $\gamma t=10$, for the initial condition $(l,m)=(4,\pm 1)$, under the condition of a fixed interface radius (no BP effects).
Figure 12 shows the second order amplitudes at the final time under the no‐BP scenario. While the shape of the spectrum is similar to that in Figure 5 (c), the magnitudes of the amplitudes are roughly three orders of magnitude smaller. This observation indicates that the BP effects exerts an even stronger influence on the growth of higher order mode couplings than it does on first order growth, underscoring the importance of incorporating BP effects in our analysis.
V Discussions and Conclusions
In this work, we developed a weakly nonlinear, multi-mode model to investigate the Rayleigh–Taylor instability (RTI) on a dynamic, three-dimensional spherical interface applicable to arbitrary initial conditons. Our model offers a comprehensive framework for analyzing RTI on a time-varying spherical surface, moving beyond many of the simplifying assumptions used in previous studies. Our analysis demonstrates that a complete description of RTI in such geometries must account for both the Bell–Plesset effects and nonlinear mode coupling, as these phenomena fundamentally alter the perturbation dynamics. The primary conclusion of our study is the identification of a powerful selection rule governing the nonlinear energy transfer between modes. We find that axisymmetric modes (those with azimuthal index $m=0$) preferentially gain energy from other modes, making this specific class of perturbations the most susceptible to rapid, destabilizing growth. This signature is fundamentally different from that observed in RTI without BP effects, where mode growth from a single mode input is largest near $m\approx(l+1)/2$ [RTI_3d_spherical1]. The discrepancy arises because the driving mechanisms are distinct: in our dynamic case, the instability is driven by the time-dependent interface deceleration, whereas in the classic static case, it is driven by a constant gravitational field.
Moreover, our results show that for a modest compression ratio $R_{0}/R_{f}=5$, the instability amplitude can be amplified by an order of $10^{2}$. This finding aligns with the scaling arguments presented by R. Epstein [BPe3]. Importantly, we further show that this amplification becomes progressively more pronounced at higher perturbation orders. The geometric convergence therefore has an even stronger impact on the growth of second and higher order mode couplings beyond the first order growth.
The theoretical framework presented here is robust and can be extended to other scenarios beyond the idealized implosion studied. For example, it is directly applicable to explosive events, or any arbitrary interface evolution. Future investigations could also incorporate more complex physics, such as time-dependent Atwood number and $\gamma_{\rho}$, and multi-mode initial conditions derived from experimental data, to explore their influence on the mode coupling dynamics. Expanding the spherical harmonic truncation to higher values of $l_{\rm max}$ would allow inclusion of smaller-scale perturbations and yield more precise quantitative predictions.
Ultimately, the insights from this study have direct implications for several key scientific fields. In astrophysics, the preferential growth of low $l$, axisymmetric modes may help explain the large-scale mixing observed in core-collapse supernovae [SNe1, SNe2, SNe3, SNe4]. In ICF, our results highlight the critical importance of minimizing axisymmetric imperfections on capsule surfaces to control instability and achieve successful ignition [Hurricane1, Hurricane2, Hurricane3, Lawrence2023, Samulski2024].
VI Acknowledgments
The authors gratefully acknowledge Professor Lifeng Wang and Dr. Jing Zhang for their insightful discussions and valuable feedback. We also extend our thanks to Zhiyuan College, Shanghai Jiao Tong University for providing a stimulating research environment that facilitated our investigation into this fascinating field.
Appendix A Equations of Conservation of Mass and Momentum
We here demonstrate how Eqs. 4 and 5 are expanded up to second order. For Eq. 4, we have
$\frac{\partial\eta}{\partial t}+\frac{1}{r^{2}\sin^{2}\theta}\frac{\partial\eta}{\partial\varphi}\frac{\partial\phi^{i}}{\partial\varphi}+\frac{1}{r^{2}}\frac{\partial\eta}{\partial\theta}\frac{\partial\phi^{i}}{\partial\theta}-\frac{\partial\phi^{i}}{\partial r}+\left(\dot{R}-\frac{\partial\psi}{\partial r}\right)=0,\quad\text{at }r=R+\eta.$ (37)
The surface deformation can be expressed as
$\eta=\sum_{l,m}a_{l,m}Y_{l,m}=\sum_{l,m}\left(\epsilon a_{l,m}^{(1)}+\epsilon^{2}a_{l,m}^{(2)}\right)Y_{l,m},$ (38)
and, as an example, the internal potential $\phi^{\mathrm{in}}$ takes the form
$\phi^{\mathrm{in}}=\sum_{l,m}\left(\epsilon b_{l,m}^{\mathrm{in}(1)}+\epsilon^{2}b_{l,m}^{\mathrm{in}(2)}\right)\left(\frac{r}{R}\right)^{l}Y_{l,m}.$ (39)
Since both Eqs. 4 and 5 are evaluated at $r=R+\eta$, it is crucial to note that after differentiating $\psi$ and $\phi^{i}$ with respect to $r$, $\theta$, and $\varphi$, all terms containing $r$ must be expanded using $r=R+\eta$ in order to obtain the complete $\epsilon$-expansion. To illustrate this procedure, consider $\partial_{r}\phi^{\mathrm{in}}$ in Eq. 4:
$\begin{split}\frac{\partial\phi^{\mathrm{in}}}{\partial r}&=\sum_{l,m}\frac{l(\epsilon b_{l,m}^{\mathrm{in}(1)}+\epsilon^{2}b_{l,m}^{\mathrm{in}(2)})}{R}\left(\frac{r}{R}\right)^{l-1}Y_{l,m}=\sum_{l,m}\frac{l(\epsilon b_{l,m}^{\mathrm{in}(1)}+\epsilon^{2}b_{l,m}^{\mathrm{in}(2)})}{R}\left(1+(l-1)\frac{\eta}{R}\right)Y_{l,m}\\ &=\sum_{l,m}\Bigg\{\epsilon\!\left[\frac{l}{R}b_{l,m}^{\mathrm{in}(1)}Y_{l,m}\right]+\epsilon^{2}\!\left[\frac{l}{R}b_{l,m}^{\mathrm{in}(2)}Y_{l,m}+\frac{l(l-1)}{R^{2}}\!\sum_{l_{1},m_{1}}\!b_{l,m}^{\mathrm{in}(1)}a_{l_{1},m_{1}}^{(1)}Y_{l,m}Y_{l_{1},m_{1}}\right]\Bigg\}.\end{split}$ (40)
All other terms are expanded analogously. Two additional useful relations are
$\frac{1}{r}\simeq\frac{1}{R}\!\left[1-\frac{\eta}{R}+\frac{\eta^{2}}{R^{2}}\right],\qquad\frac{1}{r^{2}}=\frac{1}{R^{2}(1+\eta/R)^{2}}\simeq\frac{1}{R^{2}}\!\left[1-\frac{2\eta}{R}+\frac{3\eta^{2}}{R^{2}}\right],\qquad\frac{\eta}{R}\ll 1.$ (41)
Following this procedure, both Eq. 4 and Eq. 5 can be fully expanded to second order. Upon projecting the resulting expressions onto a particular spherical harmonic mode, one obtains the first- and second-order relations summarized in Eqs. 14, 15, LABEL:eq:first_order_eq, 25, 26, and 27.
Appendix B Angular Projections and Mode Coupling
To fully project Eq. 4 and Eq. 5 onto a specific spherical harmonic mode, we first multiply both equations by $\sin^{2}\theta$. The $\sin^{2}\theta$ term itself contains spherical harmonic components; therefore, after projection, several types of mode coupling arise. In this section, we define the corresponding coefficients $A_{l,m,l_{1},m_{1}}$, $B_{l,m,l_{1},m_{1},l_{2},m_{2}}$, $H_{l,m,l_{1},m_{1},l_{2},m_{2}}$, and $\Theta_{l,m,l_{1},m_{1},l_{2},m_{2}}$, which quantify the coupling among different angular modes.
Before giving the explicit forms of the coefficients, we simplify the mode coupling integrals as they typically take the form
$\sum_{l_{1},m_{1},l_{2},m_{2}}\int Y_{l,m}(\hat{\mathbf{r}})Y_{l_{1},m_{1}}(\hat{\mathbf{r}})Y_{l_{2},m_{2}}(\hat{\mathbf{r}})\,\mathrm{d}\Omega,$ (42)
which describe how initially independent spherical harmonic modes interact. Even a single-mode perturbation can excite additional modes at second order. These integrals can be conveniently expressed using the Wigner 3jj-symbols [Wigner1993, Aquilanti2007]. The product of two spherical harmonics can be expanded as
[…]
where the asterisk denotes complex conjugation and $\begin{pmatrix}l&l_{1}&l_{2}\\ m&m_{1}&m_{2}\end{pmatrix}$ is the Wigner 3jj-symbol. Integrating over the solid angle gives the well-known identity
[…]
The 3jj-symbols vanish unless the following selection rules are satisfied:
$|l_{1}-l_{2}|\leq l\leq l_{1}+l_{2},\quad m=m_{1}+m_{2}.$ (45)
This property greatly simplifies computation and clarifies the physical origin of multi-mode interactions.
Since
$\sin^{2}\theta=\frac{2}{3}-\frac{4\sqrt{5\pi}}{15}Y_{2,0},$ (46)
the first term describes the projection of $Y_{2,0}$ onto $Y_{l,m}$. We define
$A_{l,m,l_{1},m_{1}}$ $=-\frac{4\sqrt{5\pi}}{15}\int Y_{l,m}^{*}Y_{l_{1},m_{1}}Y_{2,0}\sin\theta\,\mathrm{d}\theta\,\mathrm{d}\varphi$ (47)
$=-\frac{4\sqrt{5\pi}}{15}(-1)^{m}\int Y_{l,-m}Y_{l_{1},m_{1}}Y_{2,0}\sin\theta\,\mathrm{d}\theta\,\mathrm{d}\varphi$
$=-\frac{4\sqrt{5\pi}}{15}(-1)^{m}\sqrt{\frac{(2l+1)(2l_{1}+1)\times 5}{4\pi}}\begin{pmatrix}l&l_{1}&2\\ 0&0&0\end{pmatrix}\begin{pmatrix}l&l_{1}&2\\ -m&m_{1}&0\end{pmatrix}$,
which is nonzero only when $|l_{1}-2|\leq l\leq l_{1}+2$ and $m_{1}=m$.
The $\frac{2}{3}$ factor implies that a projection of $\frac{2}{3}Y_{l_{1},m_{1}}Y_{l_{2},m_{2}}$ contributes to the target mode $Y_{l,m}$. Thus, we define
$B_{l,m,l_{1},m_{1},l_{2},m_{2}}$ $=\int Y_{l,m}^{*}(\hat{\mathbf{r}})Y_{l_{1},m_{1}}(\hat{\mathbf{r}})Y_{l_{2},m_{2}}(\hat{\mathbf{r}})\,\mathrm{d}\Omega$ (48)
$=(-1)^{m}\sqrt{\frac{(2l+1)(2l_{1}+1)(2l_{2}+1)}{4\pi}}\begin{pmatrix}l&l_{1}&l_{2}\\ 0&0&0\end{pmatrix}\begin{pmatrix}l&l_{1}&l_{2}\\ -m&m_{1}&m_{2}\end{pmatrix}$.
Additionally, by interacting with $\frac{4\sqrt{5\pi}}{15}Y_{2,0}$, another projection term arises,
$C_{l,m,l_{1},m_{1},l_{2},m_{2}}$ $=-\frac{4\sqrt{5\pi}}{15}(-1)^{m}\int Y_{l,-m}Y_{l_{1},m_{1}}Y_{l_{2},m_{2}}Y_{2,0}\sin\theta\,\mathrm{d}\theta\,\mathrm{d}\varphi$ (49)
$=(-1)^{m}\sum_{l_{3}}A_{l_{3},m_{3},l_{2},m_{2}}\int Y_{l,-m}Y_{l_{1},m_{1}}Y_{l_{3},m_{3}}\sin\theta\,\mathrm{d}\theta\,\mathrm{d}\varphi$
$=\sum_{l_{3}}A_{l_{3},m_{3},l_{2},m_{2}}\,B_{l,m,l_{1},m_{1},l_{3},m_{3}},$
where the intermediate mode $Y_{l_{3},m_{3}}$ serves as a coupling bridge between $Y_{l_{2},m_{2}}$ and $Y_{2,0}$. This recursive formulation allows efficient use of Wigner 3jj algebra to simplify higher order products.
The total projection coefficient for $\sin^{2}\theta$ acting on $Y_{l_{1},m_{1}}Y_{l_{2},m_{2}}$ is then
$H_{l,m,l_{1},m_{1},l_{2},m_{2}}=\frac{2}{3}B_{l,m,l_{1},m_{1},l_{2},m_{2}}+C_{l,m,l_{1},m_{1},l_{2},m_{2}}.$ (50)
The last type of projection to consider involves the term $\partial_{\theta}Y_{l_{1},m_{1}}\,\partial_{\theta}Y_{l_{2},m_{2}}\,\sin^{2}\theta$, which must be projected onto the target spherical harmonic mode $Y_{l,m}$. This projection defines the coefficient $\Theta_{l,m,l_{1},m_{1},l_{2},m_{2}}$. To start with, we have
$\frac{\partial Y_{l,m}}{\partial\varphi}=imY_{l,m},\qquad\frac{\partial Y_{l,m}}{\partial\theta}=\frac{1}{2}e^{-i\varphi}G_{1}(l,m)\,Y_{l,m+1}-\frac{1}{2}e^{i\varphi}G_{2}(l,m)\,Y_{l,m-1},$ (51)
where we have defined
$G_{1}(l,m)=\sqrt{(l-m)(l+m+1)},\qquad G_{2}(l,m)=\sqrt{(l+m)(l-m+1)}.$ (52)
We then define
$\Theta_{l,m,l_{1},m_{1},l_{2},m_{2}}$ $=\int Y_{l,m}^{*}\,\sin^{2}\theta\,\bigg[\frac{1}{2}e^{-i\varphi}G_{1}(l_{1},m_{1})\,Y_{l_{1},m_{1}+1}-\frac{1}{2}e^{i\varphi}G_{2}(l_{1},m_{1})\,Y_{l_{1},m_{1}-1}\bigg]$ (53)
$\quad\times\bigg[\frac{1}{2}e^{-i\varphi}G_{1}(l_{2},m_{2})\,Y_{l_{2},m_{2}+1}-\frac{1}{2}e^{i\varphi}G_{2}(l_{2},m_{2})\,Y_{l_{2},m_{2}-1}\bigg]\sin\theta\,\mathrm{d}\theta\,\mathrm{d}\varphi,$
which expands to
$\Theta_{l,m,l_{1},m_{1},l_{2},m_{2}}$ $=\frac{1}{4}\int Y_{l,m}^{*}\,\Bigg\{$ (54)
$\quad\sin^{2}\theta\,e^{-2i\varphi}\,G_{1}(l_{1},m_{1})G_{1}(l_{2},m_{2})\,Y_{l_{1},m_{1}+1}Y_{l_{2},m_{2}+1}$
$\quad-\sin^{2}\theta\,G_{1}(l_{1},m_{1})G_{2}(l_{2},m_{2})\,Y_{l_{1},m_{1}+1}Y_{l_{2},m_{2}-1}$
$\quad-\sin^{2}\theta\,G_{2}(l_{1},m_{1})G_{1}(l_{2},m_{2})\,Y_{l_{1},m_{1}-1}Y_{l_{2},m_{2}+1}$
$\quad+\sin^{2}\theta\,e^{2i\varphi}\,G_{2}(l_{1},m_{1})G_{2}(l_{2},m_{2})\,Y_{l_{1},m_{1}-1}Y_{l_{2},m_{2}-1}\Bigg\}\sin\theta\,\mathrm{d}\theta\,\mathrm{d}\varphi.$
Using the relations
$\sin^{2}\theta$ $=\frac{2}{3}-\frac{4\sqrt{5\pi}}{15}Y_{2,0}$ (55)
$Y_{2,2}$ $=\frac{1}{4}\sqrt{\frac{15}{2\pi}}\sin^{2}\theta e^{2i\varphi}\Rightarrow e^{2i\varphi}\sin^{2}\theta=4\sqrt{\frac{2\pi}{15}}Y_{2,2},$
$Y_{2,-2}$ $=\frac{1}{4}\sqrt{\frac{15}{2\pi}}\sin^{2}\theta e^{-2i\varphi}\Rightarrow e^{-2i\varphi}\sin^{2}\theta=4\sqrt{\frac{2\pi}{15}}Y_{2,-2},$
we finally obtain
$\begin{gathered}\Theta_{l,m,l_{1},m_{1},l_{2},m_{2}}=-\frac{1}{4}G_{1}(l_{1},m_{1})G_{2}(l_{2},m_{2})H_{l,m,l_{1},m_{1}+1,l_{2},m_{2}-1}\\ -\frac{1}{4}G_{1}(l_{2},m_{2})G_{2}(l_{1},m_{1})H_{l,m,l_{1},m_{1}-1,l_{2},m_{2}+1}\\ \\ +\sqrt{\frac{2\pi}{15}}G_{2}(l_{1},m_{1})G_{2}(l_{2},m_{2})D_{l,m,l_{1},m_{1}-1,l_{2},m_{2}-1}\\ +\sqrt{\frac{2\pi}{15}}G_{1}(l_{1},m_{1})G_{1}(l_{2},m_{2})E_{l,m,l_{1},m_{1}+1,l_{2},m_{2}+1},\end{gathered}$ (56)
where
$D_{l,m,l_{1},m_{1},l_{2},m_{2}}$ $=\int Y_{l,m}^{*}Y_{2,2}Y_{l_{1},m_{1}}Y_{l_{2},m_{2}}\sin\theta\,\mathrm{d}\theta\,\mathrm{d}\varphi$ (57)
$=\sum_{l_{3}}B_{l_{3},m_{3},l_{2},m_{2},2,2}\,B_{l,m,l_{1},m_{1},l_{3},m_{3}},$
and
$E_{l,m,l_{1},m_{1},l_{2},m_{2}}$ $=\int Y_{l,m}^{*}Y_{2,-2}Y_{l_{1},m_{1}}Y_{l_{2},m_{2}}\sin\theta\,\mathrm{d}\theta\,\mathrm{d}\varphi$ (58)
$=\sum_{l_{3}}B_{l_{3},m_{3},l_{2},m_{2},2,-2}\,B_{l,m,l_{1},m_{1},l_{3},m_{3}}.$
These definitions enable the systematic decomposition of all angular products and derivative couplings in terms of Wigner 3jj-symbols, providing both analytical clarity and computational efficiency.
Appendix C Matrix Form of Second Order Equations
Following the index mapping 30, the matrix $U$ takes the form
$U(k,k_{1})=\frac{2}{3}\delta_{k}^{k_{1}}+A_{k,k_{1}},$ (59)
where $\delta_{x}^{y}$ denotes the Kronecker delta function, and $A_{k,k_{1}}$ represents $A_{l,m,l_{1},m_{1}}$ for the sets of $(l_{i},m_{i})$ corresponding to $k_{i}$. We adopt this convention throughout the remaining derivations. Under this notation, the matrices $V_{i}$ and the vectors $Q_{i}$ and $P_{i}$ are expressed as
$V_{\text{in}}(k,k_{1})=-\frac{2}{3}\frac{l}{R}\delta_{k}^{k_{1}}-\frac{l_{1}}{R}A_{k,k_{1}},$ (60)
$V_{\text{ex}}(k,k_{1})=\frac{2}{3}\frac{l+1}{R}\delta_{k}^{k_{1}}+\frac{l_{1}+1}{R}A_{k,k_{1}},$ (61)
$Q_{\text{in}}(k)=\frac{1}{R^{2}}\sum_{k_{1},k_{2}}\Big\{-m_{1}m_{2}a^{(1)}_{k_{1}}b^{\text{in}(1)}_{k_{2}}B_{k,k_{1},k_{2}}-\big[(3\dot{R}+R\gamma_{\rho})a^{(1)}_{k_{1}}a^{(1)}_{k_{2}}+l_{2}(l_{2}-1)a^{(1)}_{k_{1}}b^{\text{in}(1)}_{k_{2}}\big]H_{k,k_{1},k_{2}}+a^{(1)}_{k_{1}}b^{\text{in}(1)}_{k_{2}}\Theta_{k,k_{1},k_{2}}\Big\},$ (62)
$Q_{\text{ex}}(k)=\frac{1}{R^{2}}\sum_{k_{1},k_{2}}\Big\{-m_{1}m_{2}a^{(1)}_{k_{1}}b^{\text{ex}(1)}_{k_{2}}B_{k,k_{1},k_{2}}-\big[(3\dot{R}+R\gamma_{\rho})a^{(1)}_{k_{1}}a^{(1)}_{k_{2}}+(l_{2}+1)(l_{2}+2)a^{(1)}_{k_{1}}b^{\text{ex}(1)}_{k_{2}}\big]H_{k,k_{1},k_{2}}+a^{(1)}_{k_{1}}b^{\text{ex}(1)}_{k_{2}}\Theta_{k,k_{1},k_{2}}\Big\},$ (63)
$\begin{split}P_{\text{in}}(k)=\sum_{k_{1},k_{2}}\Bigg\{&\frac{l_{2}}{R}a^{(1)}_{k_{1}}\dot{b}^{\text{in}(1)}_{k_{2}}-l_{2}^{2}\frac{\gamma_{R}}{R}a^{(1)}_{k_{1}}b^{\text{in}(1)}_{k_{2}}+\frac{1}{R^{2}}\bigg(\dot{R}^{2}-R\ddot{R}-\frac{\dot{\gamma}_{\rho}R^{2}}{2}\bigg)a^{(1)}_{k_{1}}a^{(1)}_{k_{2}}\\ &+\frac{1}{R}\Big[l_{2}(l_{2}-1)\gamma_{R}-l_{2}(2\gamma_{R}+\gamma_{\rho})\Big]a^{(1)}_{k_{1}}b^{\text{in}(1)}_{k_{2}}+\frac{(2\gamma_{R}+\gamma_{\rho})^{2}}{2}a^{(1)}_{k_{1}}a^{(1)}_{k_{2}}+\frac{l_{1}l_{2}}{2R^{2}}b^{\text{in}(1)}_{k_{1}}b^{\text{in}(1)}_{k_{2}}\Bigg\}H_{k,k_{1},k_{2}}\\ &+\frac{1}{2R^{2}}\sum_{k_{1},k_{2}}\bigg({b^{\text{in}(1)}_{k_{1}}b^{\text{in}(1)}_{k_{2}}}\Theta_{k,k_{1},k_{2}}-m_{1}m_{2}b^{\text{in}(1)}_{k_{1}}b^{\text{in}(1)}_{k_{2}}B_{k,k_{1},k_{2}}\bigg),\end{split}$ (64)
$\begin{split}P_{\text{ex}}(k)=\sum_{k_{1},k_{2}}\Bigg\{&-(l_{2}+1)^{2}\frac{\gamma_{R}}{R}a^{(1)}_{k_{1}}b^{\text{ex}(1)}_{k_{2}}-\frac{l_{2}+1}{R}a^{(1)}_{k_{1}}\dot{b}^{\text{ex}(1)}_{k_{2}}+\frac{1}{R^{2}}\bigg(\dot{R}^{2}-R\ddot{R}-\frac{\dot{\gamma}_{\rho}R^{2}}{2}\bigg)a^{(1)}_{k_{1}}a^{(1)}_{k_{2}}\\ &+\frac{1}{R}\Big[(l_{2}+1)(l_{2}+2)\gamma_{R}+(l_{2}+1)(2\gamma_{R}+\gamma_{\rho})\Big]a^{(1)}_{k_{1}}b^{\text{ex}(1)}_{k_{2}}\\ &+\frac{(2\gamma_{R}+\gamma_{\rho})^{2}}{2}a^{(1)}_{k_{1}}a^{(1)}_{k_{2}}+\frac{(l_{1}+1)(l_{2}+1)}{2R^{2}}b^{\text{ex}(1)}_{k_{1}}b^{\text{ex}(1)}_{k_{2}}\Bigg\}H_{k,k_{1},k_{2}}\\ &+\frac{1}{2R^{2}}\sum_{k_{1},k_{2}}\bigg(b^{\text{ex}(1)}_{k_{1}}b^{\text{ex}(1)}_{k_{2}}\Theta_{k,k_{1},k_{2}}-m_{1}m_{2}b^{\text{ex}(1)}_{k_{1}}b^{\text{ex}(1)}_{k_{2}}B_{k,k_{1},k_{2}}\bigg).\end{split}$ (65)
Here, $B_{k,k_{1},k_{2}}$, $H_{k,k_{1},k_{2}}$, and $\Theta_{k,k_{1},k_{2}}$ respectively represent $B_{l,m,l_{1},m_{1},l_{2},m_{2}}$, $H_{l,m,l_{1},m_{1},l_{2},m_{2}}$, and $\Theta_{l,m,l_{1},m_{1},l_{2},m_{2}}$ under the same mapping between $(l_{i},m_{i})$ and $k_{i}$.
Appendix D Convergence Test of Truncation Index $l_{\mathrm{max}}$
[…]
To determine an appropriate truncation for the spherical harmonic expansion, $l_{\text{max}}$, we perform numerical tests using the $(l,m)=(4,\pm 1)$ mode as the initial condition and vary $l_{\text{max}}$. Figure 13 shows the second order amplitudes at the end of the evolution for several representative modes. The $l=2$ and $l=4$ modes converge rapidly at $l_{\text{max}}\leq 12$, while higher-$l$ modes converge more slowly due to their proximity to the truncation limit. Overall, the low-$l$ modes that are of interests already converge at $l_{\text{max}}=16$.